) Suppose the production function is q = 12 K.5L.5

. Calculate the MRTS (MPK/MPL) when: a) K = 200, L = 300 b) K = 50, L = 400 c) K= 900, L = 2000
Answer #1

MRTS=L/K

1.MRTS=300/200=2/3

2.MRTS=400/50=8

3.MRTS=2000/900=2.23